OpenStudy (anonymous):

I still dont understand the whole algebra stuff number and letters together confuse me. ? 3n-5=7n+11?

OpenStudy (nincompoop):

put n to the left side 3n-7n -5 + 5n = 7n -7n + 11 + 5n

OpenStudy (anonymous):

3n-5=7n+11 3n - 7n = 11+ 5 -4n = 16 -n = 16/4 -n = 4 n= -4